OpenStudy (anonymous):

Solve the system for the x-coordinate of the point of intersection. 3x + y = -2 2x - 5y = 4

OpenStudy (anonymous):

Can Someone Please Give me The Answer For this problem?

sam (.sam.):

3x + y = -2 2x - 5y = 4 -------------------------------------------------------------------- 3x + y = -2 y =-3x -2 -------------------------------------- 2x - 5y = 4 2x-5(-3x-2)=4 2x+15x+10=4 17x+10=4 17x=-6 x=-(6)/(17)

OpenStudy (anonymous):

Solve the system of equations for the point of intersection of the two lines. 3x - 4y = -20 y = -3x -10